### v1

# BoostStep: Boosting mathematical capability of Large Language Models via improved single-step reasoning

Abstract
Cutting-edge large language models (LLMs) demonstrate promising performance in solving complex math problems with a divide-and-conquer pipeline and the assistance of in-context learning (ICL) examples. However, their potential for improvement is limited by two critical problems within their ICL examples: granularity-mismatch and the ensuing negative-effect noise problem. Specifically, the LLMs are capable of the dividing process yet mostly failed by inaccurate reasoning within a few conquer steps, while the ICL examples retrieved in question-grained sometimes lack relevant steps for a specific challenging reasoning step. Further, this disconnect may hinder the correct reasoning due to its irrelevance. To this end, we focus on improving the reasoning quality within each step and present BoostStep. BoostStep aligns the granularity between the retrieving and reasoning on step grained, and provides highly related ICL examples for each reasoning step with a novel ‘first-try’ strategy. BoostStep provides more relevant examples than the coarse question-grained strategy, enhancing the model reasoning quality within each step steadily. BoostStep is a general and robust reasoning-enhancing method that not only improves standalone reasoning performance but also integrates seamlessly with Monte Carlo Tree Search methods (MCTS) to refine both candidate generation and decision-making. Quantitatively, it improves GPT-4o and Qwen2.5-Math-72B by 3.6% and 2.0% respectively on various mathematical benchmarks, and 7.5% gain combined with MCTS.
BoostStep: Boosting mathematical capability of Large Language Models via improved single-step reasoning
Beichen Zhang∗1,2, Yuhong Liu∗1,2, Xiaoyi Dong†1,3, Yuhang Zang1, Pan Zhang1, Haodong Duan1, Yuhang Cao1, Dahua Lin1,3, Jiaqi Wang†1 1Shanghai AI Laboratory  2Shanghai Jiao Tong University 3The Chinese University of Hong Kong https://github.com/beichenzbc/BoostStep
footnotetext: * indicates equal contribution $\dagger$  indicates corresponding author
1 Introduction
Figure 1: Our step-level in-context learning outperforms traditional problem-level few-shot learning for about 4% across in-domain, out-domain and cross-modality mathematical benchmark. Moreover, on benchmarks with lower similarity with example problem set (i.e. OlympiadBench and multi-modal benchmarks), where problem-level few-shot learning may have negative impact, while our strategy still provides valuable guidance.
Mathematical reasoning is a crucial and challenging task in the development of artificial intelligence. It serves as an indicator of a model’s ability to perform complex reasoning and has a wide range of applications, such as problem solving, theorem proving and scientific discovery.
When solving complex mathematical problems, two capabilities are necessary, which are divide and conquer. The former decomposes the complex problem into several simpler steps to simplify the task, while the latter guarantees the correctness of the one-step reasoning.
Through the analysis of error cases, we found that the current SOTA models have been relatively clear in the divide part, that is, the model can basically know exactly what tasks should be completed in each step. However, there will still be a lot of mistakes in the conquer section, such as wrong formula use, wrong calculation, insufficient enumeration, etc. To quantitatively substantiate this conclusion, we provide GPT-4o-mini with groundtruth reasoning process to determine whether the error in another response were due to an overarching flawed reasoning approach or a deviation within a particular step. In less advanced models like LLaMA-3.1-8B-Instruct Dubey et al. (2024), 8.7% of errors originates from an overall flawed reasoning approach. While in more advanced models like GPT-4o, only 0.8% of errors are ascribable to the overall approach, with the majority resulting from inaccuracies at individual steps of the reasoning process. Therefore, it can be concluded that the correctness of one-step reasoning is the bottleneck of reasoning capability.
Various approaches have been employed to improve reasoning correctness, such as producing chains of thought through prompt engineering Kojima et al. (2022); Wei et al. (2022), fine-tuning with mathematical data Shao et al. (2024); Yang et al. (2024); Ying et al. (2024), or generating multiple candidate reasoning paths using Monte Carlo Tree Search (MCTS) Zhang et al. (2024b, a).
Among those techniques, in-context learning is a particularly important one, which offers similar examples to provide detailed guidance. However, the examples retrieved by traditional problem-level in-context learning are listed before the reasoning process ever begins, thereby lacking fine-grained guidance during the reasoning process. Moreover, since the example problem can’t be exactly identical to the new one, the irrelevant steps in those examples may even become a distraction from the current reasoning, thus even negatively affecting the conquer capability for some specific steps.
To this end, we refine in-context learning from problem-level to step-level granularity to offer similar example steps during an on-going reasoning process for fine-grained step-level guidance, as well as ensuring that the introduced example is still highly relevant at the step level to avoid distractions.
Firstly, we have constructed an example problem bank with step-level granularity based on reasoning content, instead of commonly adopted grammatical separation. This ensures the steps in the question bank are consistent with the actual reasoning steps, thereby providing more appropriate guidance.
Building on the step-level granularity within the example problem bank, we propose an approach that incorporates in-context learning through a "first-try" format during an on-going reasoning process. Specifically, for a given problem to be solved, we break down the solving process into step-by-step reasoning paths. During the reasoning of a single step, we first allow the model to attempt a "first try" to comprehend what the model currently needs to reason about. Based on this initial attempt, we search the question bank to find similar steps that can guide the model to accurately output the current step. This first-try strategy helps ensure a higher similarity between the retrieved examples and the current step. Therefore, the distraction from irrelevant steps can be avoided and and the guidance effect is improved.
Compared with traditional in-context learning in problem-level granularity, our method is capable of providing examples during the reasoning process directly based on the steps to be solved, thereby offering more relevant guidance. Our approach demonstrates significant improvements over traditional few-shot learning across various open-source benchmarks, with an average increase of 4%.
Moreover, our method also reduces the dependency on the similarity between the example dataset and the target problem since two entirely different problems can still share similar key steps, such as utilizing prime factorization for both calculating the least common multiple and determining divisibility. Consequently, dissimilar problems can still offer effective guidance. For example, on multi-modal benchmarks with lower similarity to the example problem bank like MathVision Wang et al. (2024), traditional few-shot learning can even have negative effects. In contrast, our approach continues to achieve significant improvements.
Our method significantly enhances single-step reasoning accuracy through step-level in-context learning, making it highly compatible with various current reasoning strategies that employ Monte Carlo Tree Search (MCTS). Typically, an MCTS method requires a reasoning model to generate multiple step-wise candidate reasoning paths and a critic model to evaluate the correctness of these candidates. Our approach can be integrated into both aspects in a plug-and-play manner. Specifically, when generating new candidate reasoning nodes, our method can introduce similar examples in the aforementioned ’first-try’ manner to improve the accuracy of candidates. Additionally, it can aid the critic model by incorporating similar correct reasoning steps into the evaluation of candidate reasoning processes. Experimental results indicate that both applications contribute positively to the model’s reasoning capabilities during MCTS process. Detailed illustration will be shown in Sec. 3.4
2 Related Works
Mathematical Reasoning. Mathematical reasoning has long been a highly challenging task in the field of artificial intelligence. In the early days of artificial intelligence, constrained by a lack of general capabilities, early methods Feigenbaum et al. (1963); Fletcher (1985) primarily attempted to perform simple mathematical reasoning through rule-based methods. With the advent of large language models with enhanced reasoning capabilities, contemporary approaches typically focus on enhancing performance during both the training and inference phases. The first category improves mathematical capability by fine-tuning with more high-quality mathematical data Shao et al. (2024); Yang et al. (2024); Lewkowycz et al. (2022); Yue et al. (2023); Xu et al. (2024). This strategy can fundamentally improve the base model’s mathematical capabilities. However, it demands substantial high-quality mathematical data and computational resources. Consequently, more efforts have been put in exploring various techniques during inference to enhance mathematical reasoning performance. Some work Wei et al. (2022); Kojima et al. (2022) involves prompt engineering to enable models to generate comprehensive chains of thought. Other studies Madaan et al. (2024); Gou et al. (2023); Ke et al. (2024) involve using self-refinement techniques to revise the initial reasoning outputs.
Stepwise Mathematical Reasoning. Recently, to further enhance mathematical reasoning capabilities, many studies have shifted the granularity of mathematical reasoning from the problem level to the step level. This approach involves addressing each next step individually, completing small segments of reasoning within the overall task. These works often employ strategies such as Tree of Thoughts (ToT) Yao et al. (2024) or Monte Carlo Tree Search Zhang et al. (2024b, a); Chen et al. (2024); Feng et al. (2023); Zhu et al. (2022), extending multiple steps to optimize step answers and ultimately obtain the optimal solution. Additionally, Process Supervision Models (PRMs) Lightman et al. (2023); Luo et al. (2024) are frequently used to verify the correctness of new candidate nodes in real-time and prune reasoning paths, thereby improving the accuracy of the final answer. This more detailed auxiliary strategy demonstrates greater potential.
In-context Learning in Mathematical Reasoning.  In-context learning can provide low-cost guidance to models through similar examples, thereby enhancing the quality of model outputs and their ability to follow prompts. Consequently, it has been widely adopted. However, research on in-context learning within mathematical reasoning tasks remains insufficient. Typically, this approach involves providing the model with similar problems and their ground truth solutions to offer a general strategy for solving new problems Hendrycks et al. (2021); Wei et al. (2022). Some efforts have been made to improve the relevance of retrieved examples by designing better retrieval mechanisms and incorporating appropriate reference rejection techniques Liu et al. (2024b). Others try to provide high-level context instead to improve the generalizability Wu et al. (2024). However, all these methods share a common limitation: the lack of fine-grained real-time guidance at the step level.
3 Step-Level In-Context Learning
Figure 2: Our strategy refines in-context learning from problem-level granularity (fig.a) to step-level granularity(fig.b) to provide more real-time fine-grained guidance. Moreover, our strategy can guide the reasoning and verifying process in Monte Carlo Tree Search (MCTS) strategies by introducing examples.
3.1 Revisiting In-Context Learning from Conditional Probability
Current models often employ next-token prediction for training and inference, where the conditional probability is central to the model’s generation of the next token. Given a problem $q$, a model’s reasoning process can be represented by $r_{predict}=\arg\max\limits_{r}P_{model}(r\mid q)$, where we train the model to get a better conditional probability $P_{model}$ so that $r_{predict}$ can be closer to the groundtruth answer $r_{gt}=\arg\max\limits_{r}P_{gt}(r\mid q)$.
In-context learning provides the model with conditional probabilities similar to the groundtruth answer for imitation without changing the probability model $P_{model}$. Specifically, an example problem $q^{\prime}$ and its corresponding correct solution $r^{\prime}$ is provided and it can be posited that the conditional probability $P(r^{\prime}\mid q^{\prime})$ is similar to the probability of the grountruth answer of the target problem $P(r_{gt}\mid q)$. Consequently, the model will imitate this similar example and $r^{\prime}_{predict}=\arg\max\limits_{r}P_{model}(r\mid q,q^{\prime},r^{\prime})$ will be closer to $r_{gt}$ comparing to $r_{predict}$.
However, given that the actual reasoning process $r$ can be highly complex, the complete reasoning process is often divided into multiple steps $s_{1},s_{2},\ldots$. Step-level reasoning iteratively guides the model to generate the next step $s^{0-shot}_{i+1}=\arg\max\limits_{s}P_{model}(s\mid q,s_{1},s_{2},\ldots,s_{i})$.
At the step granularity, examples retrieved based on the problem $q$ are evidently insufficient for providing appropriate guidance. Similar problem $q^{\prime}$ may not necessarily contain the corresponding steps to guide the reasoning for the new problem $q$. Moreover, irrelevant steps may provide dissimilar conditional probabilities, thereby distracting the model’s reasoning process.
To this end, we propose step-level in-context learning and first-try strategy to provide detailed and relevant example step when in step-level reasoning. Specifically, when generating new steps $s_{i+1}$ based on previous reasoning steps $s_{i},s_{i-1},\ldots,s_{1}$ and question $q$, we first utilize a first-try strategy to obtain an approximate estimate of $s^{first}_{i+1}$. Then, we use this $s^{first}_{i+1}$ to retrieve a similar step $s^{\prime}_{n+1}$ along with the corresponding $q^{\prime},s^{\prime}_{1},s^{\prime}_{2},\ldots,s^{\prime}_{n}$. Since these two steps are similar, a very reasonable assumption is that $P(s^{\prime}_{n+1}\mid q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n})$ closely approximates $P(s_{{gt}_{i+1}}\mid q,s_{1},\ldots,s_{i})$. Therefore, the generated step $s_{i+1}=\arg\max\limits_{s}P_{model}(s\mid q,s_{1},\ldots,s_{i},s^{\prime}_{n+1},s^{\prime}_{n},\ldots,s^{\prime}_{1},q^{\prime})$ will be more closed to $s_{{gt}_{i+1}}$ comparing to $s^{0-shot}_{i+1}$. Details about our step-level in-context learning and first-try strategy will be explained in Sec. 3.3
3.2 Step-Level Example Problem Bank
Figure 3: An example of different problems may contain similar steps. Problem-level in-context learning will ignore this example due to low problem similarity, while our step-level in-context learning can introduce the core skills by step-level retrieval and guidance.
Due to the need for further improvement in mathematical capabilities, currently open-source mathematical data no longer consist solely of problems and their final answers to determine whether the final answer obtained is correct or not. Instead, they also provide detailed solution processes to provide more fine-grained measurement. However, most current open-source mathematical data still do not break down the solution processes to the step level.
Some approaches Lightman et al. (2023) proposed using the period ’.’ as a clear semantic delimiter to segment steps for training PRM. This strategy allows for the quick decomposition of each step from a complete process without any additional assistance. However, this simple decomposition mode is obviously unreliable. Essentially, a single reasoning step should have a consistent target and encompass a complete thought process, making it the atomic granularity of reasoning. Using a period ’.’ as a delimiter may disrupt this atomicity. For example, it may split a complete enumerations for the same objective into multiple steps.
Therefore, we suggest that the most appropriate method for step segmentation is to allow the reasoning model itself to autonomously decompose the process. This approach ensures that the granularity of the decomposed steps in example problem bank aligns with that of the real-time reasoning steps. Specifically, we define the concept of a step through prompts, which encapsulate a complete and simple inference. This guides GPT-4o in decomposing the answer at the step level.
A major advantage of decomposing the question example bank into individual steps is that it facilitates step-level retrieval and guidance, which is of significant importance. As illustrated in Fig. 3, two distinctly different problems may contain similar key steps. Traditional problem-level in-context learning often overlooks such examples, whereas step-level in-context learning can effectively recall these steps, thereby providing fine-grained guidance to the ongoing reasoning process.
The proposed step-level example problem bank is available at https://github.com/beichenzbc/BoostStep
3.3 Step-Level ICL with First-try Strategy
The core challenge of in-context learning lies in how to effectively retrieve relevant problems or steps for effective guidance. This is contingent upon both the similarity between the problem database and the target problem, as well as the specific retrieval strategy employed. Traditional problem-level in-context learning involves retrieving similar problems based solely on the problem statement. This approach is relatively straightforward but effective, as similar problems typically encompass similar reasoning processes.
At the more granular step level, however, the situation becomes much more complex. A simple strategy is to perform retrieval using the given problem and all preceding reasoning steps $s_{i-1},s_{i-2},\ldots,s_{1},q$. The clear drawback of this method is the excessive length of the retrieval content, which diminishes the emphasis on the uniqueness of the current step. Another strategy is to use the previous step $s_{i-1}$ to retrieve $s^{\prime}_{j-1}$ from a step-level database, thereby guiding the reasoning of $s_{i}$ through the correct resolution of $s^{\prime}_{j}$. However, this approach is rather crude, as it models stepwise reasoning as a Markov process, which is evidently unreasonable. Similar steps can be applicable to different reasoning tasks, and therefore similarity in the previous step does not necessarily indicate that the retrieved subsequent step will provide valuable guidance for the reasoning in the current step.
To this end, we propose a straightforward and effective "first-try" strategy to enhance the similarity of search steps. Our premise is that the most accurate way to estimate the next step is to actually allow the model to attempt the reasoning for the next step. Specifically, given a problem $q$ and all preceding reasoning steps $s_{i-1},s_{i-2},\ldots,s_{1}$, we first instruct the model to attempt continuing the reasoning process to arrive at a tentative step $s_{i}^{try}$ without the aid of any examples. Subsequently, we use $s_{i}^{try}$ to retrieve similar steps $s^{\prime}_{j}$ along with their corresponding problem $q^{\prime}$ and preceding steps $s^{\prime}_{1},\ldots,s^{\prime}_{j-1}$ from a step-level database. Finally, we feed the retrieved similar steps back to the model, enabling it to deduce the final step $s_{i}$. Besides, we add a widely-accepted strategy reference rejection. Specifically, If the similarity of the retrieved most similar example remains below a certain threshold, we consider that there are no sufficiently similar examples available for reference. Consequently, we do not provide any examples in order to avoid the negative effects associated with incoherent in-context learning. This "try-retrieve-reason" strategy significantly enhances retrieval relevance, thereby improving reasoning effectiveness. Experiments in Sec. 4.4 compare our method with several other retrieval strategies, demonstrating the superiority of our approach.
3.4 Step-Level Example Guidance in MCTS
The core aspect of our method is that our step-level in-context learning can significantly enhance the model’s single-step reasoning capability, which makes it easily integrable into common step-reasoning strategies, such as Monte Carlo Tree Search (MCTS).
Generally, MCTS methods necessitate two key components: a reasoning model that generates step-level reasoning and a Process-Supervision Reward Model (PRM) that continuously evaluates the current reasoning step in real time. Our method is beneficial for both of these components. It enhances the step-level reasoning performed by the reasoning model and improves the effectiveness of the PRM in evaluating current reasoning steps.
For the reasoning model, MCTS inherently requires step-by-step reasoning expansion. When expanding at node $s_{i}$, we simply apply the previously mentioned strategy: the model performs $n$ first tries to obtain $n$ example steps. For each example, the model completes the reasoning to generate $n$ child nodes $s^{1}_{i+1},s^{2}_{i+1},\ldots,s^{n}_{i+1}$. Our strategy improves the accuracy of individual child nodes $s^{j}_{i+1}$, thereby enhancing the overall effectiveness of the MCTS strategy.
Evidently, judgment ability is closely related to reasoning ability. Therefore, since our strategy can enhance the accuracy of single-step reasoning, a reasonable assumption is that introducing appropriate example steps can similarly improve the PRM’s ability to assess the correctness of the current reasoning process. In particular, when evaluating the correctness of an inference step candidate $s^{j}_{i}$, we retrieve similar steps $s^{\prime}_{k}$ along with their corresponding preceding steps $s^{\prime}_{k-1},\ldots,s^{\prime}_{1}$ and question $q^{\prime}$ from the step-level example bank. Similarly, the probability distributions $P(s^{\prime}_{k}|s^{\prime}_{k-1},\ldots,s^{\prime}_{1},q^{\prime})$ and $P(s_{gt_{i}}|s_{i-1},\ldots,s_{1},q)$ exhibit similarities. This resemblance aids in assessing the discrepancy between $s^{j}_{i}$ and $s_{gt_{i}}$, thereby enhancing the accuracy of the critic model’s evaluations.
Detailed ablation experiments in Sec. 4.5 demonstrate that both strategies contribute positively to the overall reasoning quality of Monte Carlo Tree Search (MCTS) methods.
4 Experiments
4.1 Experiment setting
Reasoning Model.  Our primary reasoning model is GPT-4o Hurst et al. (2024). To demonstrate the generality of our approach, we also conducted tests on Qwen2.5-Math-72B-Instruct Yang et al. (2024), which is the state-of-the-art open-source model in mathematical reasoning capabilities.
Evaluation Benchmark.  We tested our approach on several challenging open-source mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a) College-level and High-level tasks. In addition, we manually collected a selection of problems from the AMC-10 and AMC-12 competitions to serve as even more challenging benchmarks. This is available at https://github.com/beichenzbc/BoostStep . To simulate benchmarks with lower similarity to the example problem bank, we also conducted tests on MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), highly challenging multi-modal math benchmarks
Retriever.  We utilized the classic TF-IDF encoding method combined with cosine similarity as the retriever for all for both problem-level and step-level in-context learning methods. The TF-IDF weight matrix is derived from the example problem bank. This is because the impact of the newly generated step on both TF and IDF is negligible, and real-time calculation of TF-IDF would require a significant amount of time.
Hyper-Parameters. The temperature value is 0 in all the experiments except for Monte Carlo Tree Search, which needs some random sampling to generate different reasoning candidates and the temperature value for MCTS is set 0.3. The reference rejection threshold is 0.7.
Prompt. Apart from some necessary guidance like step-level reasoning, we ensured that the prompts for each method were as similar as possible to make the comparison fairer. The specific prompts is listed in the supplementary materials.
4.2 Comparing to Problem-Level ICL
Table 1: A comparison of different in-context learning strategies on different benchmarks on GPT-4o and Qwen2.5-Math-72B-Instruct. The example problem bank is constructed from PRM800K, so MATH500 is an in-domain benchmark while others are all out-domain benchmarks. Best results are in bold.
Model Method in-domain out-domain
MATH AMC12 AMC10 AQUA MathBench(C) MathBench(H) OlympiadBench-TO
GPT-4o 0-shot 73.4 53.6 55.8 81.1 80.0 77.3 40.6
few-shot 73.8 56.5 56.7 83.9 80.7 79.3 39.3
Ours 76.4 63.0 60.4 85.4 82.0 84.0 43.3
Qwen 0-shot 83.0 67.4 67.7 84.6 80.6 82.0 49.7
few-shot 83.8 67.4 66.8 85.0 81.3 82.7 49.9
Ours 85.2 69.2 69.6 86.6 82.7 84.7 52.7
We conduct a rigorous comparison of the proposed step-level in-context learning approach and traditional problem-level few-shot learning across multiple benchmarks and base models. Specifically, for traditional problem-level few-shot learning, we set the shot num to 4, which is a common setting for evaluation. The results are presented in Tab. 1. Our step-level In Context-Learning achieves improvements across various mathematical benchmarks, with particularly notable enhancements observed on the relatively more challenging AMC problems.
Table 2: Comparison of different strategies in multi-modal mathematical benchmarks with lower similarity with our problem bank. Base models are all GPT-4o.
Method MathVision-Mini MathVerse-Mini
0-shot 30.6 53.2
few-shot 28.7 53.2
Ours 35.2 54.2
Table 3: Experiments on the sensitivity of the similarity between the question and the example problem bank. R_t indicates that the examples are the t_th similar for different method without any rejection strategy. Given a less similar example, our method suffers an 2.26% performance loss, which is much lower comparing to few-shot learning(4.4%).
Method Math-level5 AMC12 AMC10
0-shot 50.7 53.6 55.8
few-shot R_1 52.2 56.5 56.7
few-shot R_4 46.3 (-5.9) 52.2(-4.3) 53.7 (-3.0)
Ours R_1 56 62.3 60.4
Ours R_4 52.2 (-3.8) 61.6 (-0.7) 58.1 (-2.3)
Sinece the sensitivity to example problem banks is a critical factor limiting the generalizability of in-context learning, we have designed an additional experiment to assess the sensitivity of both step-level and problem-level in-context learning approaches. Specifically, we also test different in-context learning methods on multi-modal mathematical benchmarks including MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), which has much lower similarity with our example problem bank. The results are shown in Tab. 2. Problem-level few-shot learning not only fails to enhance reasoning performance but can also negatively impact reasoning quality, while ours continues to achieve appreciable improvements.
Moreover, we also manually decrease the similarity between the examples and the problems by selecting the t_th similar example during reasoning. The result is shown in Tab. 3. We can observe that traditional problem-level in-context learning suffers from a severe decrease and is even worse than 0-shot learning when t is larger than 4. In contrast, our method doesn’t not show a significant decline.
These experiments suggest that traditional problem-level few-shot learning requires a high similarity within the example problem bank, which limits its generalizability. Our approach, in contrast, refines the granularity to the step level. Since different problems can still contain similar steps, our method consistently achieves improvements even the examples are dissimilar, demonstrating better general applicability.
4.3 Construction of Example Problem Bank
To better align with the steps in reasoning, we propose constructing a step-level question bank based on the reasoning content rather than grammatical divisions. To prove our assumption, we compare our approach with a commonly used strategy that constructs steps based on grammatical segmentation, using periods ’.’ as the delimiter, on the same dataset PRM800K and under identical conditions. Results are presented in tab. 4. Our method largely outperforms those using periods as delimiter.
Table 4: Comparison of different step-level example problem Bank construction methods.
Strategy AMC12 AMC10 MATH
Grammatical Separation 56.5 58.1 74.8
Reasoning Content 63.0 60.4 76.4
4.4 Comparison of Retrieving Strategies
The key factor in determining the effectiveness of in-context learning lies in the relevance of the retrieved examples. At the finer-grained step level, designing an appropriate retrieval strategy becomes even more crucial and challenging. Therefore, we propose the first-try strategy, which involves understanding what the model currently needs to reason about using a first attempt, and then searching the problem set for similar steps to guide the model in fully outputting the current step. To validate the effectiveness of this method, we compare it with several other strategies mentioned in Sec.3.3, retrieving by the entire reasoning path $s_{i-1},s_{i-2},\ldots,s_{1},q$ or only by the immediately preceding step $s_{i-1}$.
Tab. 5 presents the detailed result. Our method significantly outperforms the other two retrieving strategies, better anticipating the content that needs to be inferred in the current step.
Table 5: Comparison on different retrieval strategies in step-level in-context learning. The base model is GPT-4o and all the prompts are the same. ’Path’ represents retrieving by the reasoning path including all previous step $s_{i-1},s_{i-2},\ldots,s_{1}$ and question $q$, while ’Pre-Step’ represents retrieving by only the immediately preceding step $s_{i-1}$. Best results are in bold.
Strategy AMC12 AMC10 MATH MathVision
Path 56.5 58.1 73.8 31.7
Pre-Step 57.2 56.7 74.0 31.0
First-try 63.0 60.4 76.4 35.2
4.5 Example-guided Monte Carlo Tree Search
Table 6: A detailed ablation on incorporating retrieving similar steps to provide fine-grained guidance during the reasoning and verifying phases of Monte Carlo Tree Search (MCTS) methods. Base models are GPT-4o and prompts are the same. Best results are in bold.
Reason Verify AMC12 AMC10 MATH
w/o MCTS 53.6 55.8 73.4
58.7 59.0 77.8
64.4 62.2 79.2
61.6 60.4 78.2
65.2 63.6 79.4
As discussed above, the reasoning capability of generating model and the verifying capability of critic model are two core factors of Monte Carlo Tree Search methods, and our strategy can enhance the reasoning quality of MCTS in both ways. On one hand, it can improve the accuracy of generating candidate nodes using the previously mentioned first-try strategy when reasoning nodes are generated. On the other hand, it can increase the accuracy of evaluation by introducing similar examples during critic model assessments and therefore ensures that the correct reasoning nodes are more likely to be preserved. These two strategies can be decoupled, allowing us to demonstrate the effectiveness of each component through ablation studies.
In the setup for MCTS, we utilize GPT-4o as reasoning model and GPT-4o-mini as Process-supervised Reward Model. Moreover, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b)which reduces variability and ensures a more robust evaluation. Detailed setting will be listed in appendix.
Tab. 6 presents the results of integrating in-context learning into the reasoning and evaluation phases of the Monte Carlo Tree Search methods. The results indicate that introducing examples enhances the verifying and reasoning capabilities of MCTS methods, therefore both approaches contribute to the improvement of overall reasoning performance through providing similar groundtruth steps. Specifically, introducing example steps in reasoning phases will bring more benefits to the overall reasoning quality.
4.6 Case Study
Figure 4: A specific example of adjusting reasoning during real-time inference through step-level in-context learning. The first-try uses a wrong equation while the retrieving example step guides the model to use the correct equation and get the correct conclusion.
Here we demonstrate a specific example of how our step-level in-context learning boosts step-level reasoning. Given the question, we first let the model to have a first-try on step one. Unfortunately, because the model is unfamiliar with trigonometric functions, it makes an error on tangent sum formula, therefore leading to a wrong step. However, we can get a rough idea on what the model wants to calculate at this step according to the first try. Then, we find a similar step which correctly leverage tangent sum formula in the step-level example problem bank. Therefore, with the guidance provided, the model correctly applied the tangent sum formula during the second reasoning attempt and arrived at the correct answer.
5 Conclusion
We propose step-level in-context learning, which provides real-time, fine-grained guidance during the reasoning process by searching for similar steps according to the first-try reasoning attempt. This approach improves the model’s reasoning capabilities and reduces the dependency on the similarity of example problem set, thereby increasing the generalizability of in-context learning. Moreover, our method can also enhance the reasoning and evaluation capability of Monte Carlo Tree Search (MCTS) by introducing similar steps in reasoning and verifying phases respectively, thereby improving the overall reasoning correctness.
Potential Limitations. Currently, our example problem bank is entirely sourced from PRM800k, resulting in a relatively homogeneous distribution of example problems and example steps. Furthermore, the TF-IDF retriever is based on modeling language term frequency directly and lacks an understanding of mathematical content, which limits its retrieval capabilities on math problems.
Acknowledgments
This project is funded in part by Shanghai Artificial lntelligence Laboratory, the National Key R&D Program of China (2022ZD0160201), the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. Dahua Lin is a PI of CPII under the InnoHK.
Appendix A Detailed Experiment Setting
A.1 Prompt
[…]
Example i:
Problem: xxx
Solution: xxx
[…]
Example Problem: xxx
Example Solution: Step1: xxx, Step2: xxx, …, Stepi(Key Step): xxx.
A.2 Details of Grading and Metrics
[…]
A.3 Benchmarks
We tested our approach on several challenging open-source mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a). Specifically, we use the Olympiad-TO (text-only) subset of OlympiadBench, and the application problems in college-level and high-level difficulty.
For multi-modal math benchmarks, we use MathVision-Mini Wang et al. (2024) and vision-dominant version of problems in MathVerse-Mini  Zhang et al. (2025).
Appendix B Detailed Setup for MCTS
In the setup for MCTS, we utilize GPT-4o as the reasoning model and employ GPT-4o-mini as the Process-supervised Reward Model (PRM). For the PRM, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b). Specifically, PPRM transforms the absolute rewards calculation into preference predictions between solutions to calculate rewards. This approach reduces the variability associated with scoring characteristics and thus leads to a more robust and consistent evaluation of different solutions.
The complete reasoning process of MCTS in our experiment is as follows: we start with the target problem as the root node and obtain two initial solution steps through sampling to serve as the two initial parent nodes. In each step-level reasoning phase, we expand these two parent nodes through sampling, generating four candidate child nodes. Using the PPRM, we select the two child nodes with higher confidence to become the parent nodes for the next step of reasoning. This process continues until both candidate nodes have completed their reasoning paths, resulting in the final answers. Finally, PPRM is used to select the ultimate answer from these two reasoning paths.